# Dagger Behind Smile: Fool LLMs with a Happy Ending Story

### v1

@@ Abstract
The wide adoption of Large Language Models (LLMs) has attracted significant attention from jailbreak attacks, where adversarial prompts crafted through optimization or manual design exploit LLMs to generate malicious content. However, optimization-based attacks have limited efficiency and transferability, while manual designs are either easily detectable or demand intricate interactions with LLMs. In this paper, we first point out a novel perspective for jailbreak attacks: LLMs are more responsive to positive prompts. Based on this, we deploy Happy Ending Attack (HEA) to wrap up a malicious request in a scenario template involving a positive prompt formed mainly via a happy ending, it thus fools LLMs into jailbreaking either immediately or at a follow-up malicious request. This has made HEA both efficient and effective, as it requires only up to two steps to fully jailbreak LLMs. Extensive experiments show that our HEA can successfully jailbreak on state-of-the-art LLMs, including GPT-4o, Llama3-70b, Gemini-pro, and achieves 88.79% Attack Success Rate on average. We also provide potential quantitative explanations for the success of HEA. WARNING: This paper contains offensive and harmful content generated by LLMs.
@@ 1 Introduction
Figure 1: As shown in the upper panel, an LLM can still detect maliciousness in a negative scenario camouflage, while in the lower panel, our HEA builds a positive scenario via a happy ending and fools the LLM into responding to the malicious parts in the prompt, getting jailbreak steps in the generated dialog with just one round of interaction. Content is taken from a conversation with GPT-4o.
Learning from adversarial attacks [9], many jailbreak attacks have focused on optimization-based strategies. These strategies leverage optimization algorithms to automatically refine prompts, allowing them to circumvent LLM restrictions. Gradient-based attacks such as [10, 11, 12] utilize adversarial tokens to bypass constraints on LLMs. However, these methods require access to the model’s parameters, which limits their practicality in black-box scenarios. In contrast, [13] and [14] utilize genetic algorithms to filter and refine jailbreakable prompts. Although these approaches are effective against specific LLMs, the adversarial prompts they generate suffer from limited transferability, often failing to generalize across different models. In general, methods leveraging optimization are computationally intensive and time-consuming, leading to significant loss in efficiency.
To improve the efficiency, jailbreak attacks through manual design that leverage strong logical reasoning capabilities of LLMs [15, 16] are receiving more attention [17]. For instance, DAN [18] uses simple single-round scene camouflage to jailbreak LLMs but has become easily detectable as LLMs’ safety alignment advanced [19, 20, 21]. To counter this, more sophisticated methods have been proposed including paraphrasing malicious intents into cryptic hints [22, 23, 24], decomposing a malicious question into multiple related subproblems [25, 26, 27], or employing multi-turn dialogues with extended contexts [28, 29, 30, 31, 32, 33]. However, these approaches often involve complex interactions with a target model, lack standardized templates, and suffer from instability, frequently requiring manual intervention, which limits their overall effectiveness.
Therefore, developing an effective and efficient jailbreak attack method remains a critical challenge. We observe that most existing attack templates focus on mitigating the prominence of malicious intentions within prompts to evade the security checks of LLMs. Inspired by this, We find that LLMs are more responsive to positive prompts and avoid giving respond to negative ones, which is a novel perspective for jailbreak attacks. As shown in Figure 1, if a prompt with malicious content leads to a negative impact, LLMs will basically refuse to answer the related questions; whereas, if a prompt is telling about a positive event, even though it contains a malicious request, LLMs will be able to respond to the whole prompt normally, thus inadvertently reacting to malicious parts.
Based on our findings, we propose the first jailbreak attack that exploits the positive sentimental disposition of a prompt: the Happy Ending Attack (HEA). Specifically, we embed the malicious request into a universal template that applies scenario camouflage and we give the virtual scene in the template a happy ending, making the whole story positive. This happy ending is able to fool LLMs into believing that they are giving a beneficial answer, while actually responding to the malicious request in the template as well, enabling the HEA template to get harmful outputs with only one round of dialog. To get more detailed and organized jailbreak responses, we also further design one fixed jailbreak prompt with a Chain-of-Thought (CoT) instruction based on our happy ending template to query the target LLM. HEA does not require complex conversations, but only up to these two steps. HEA also enables complete automation from template generation to conduct the attack, with effectiveness and efficiency at the same time.
We systematically tested HEA on the full AdvBench Dataset [10]. We selected metrics including attack success rate (ASR), number of tokens used for one-round attacks, and harmfulness score [34]. We tested the performance of HEA on state-of-the-art (SOTA) commercial and open-source LLMs, including GPT-4o and 4o-mini, Gemini-pro and Gemini-flash, Llama3-8b and 70b. Besides, we also provide possible explanations for the success of HEA using sentiment classification [35] and saliency heatmaps [36], which can help researchers understand the decision-making basis of LLMs better and guide the exploration of better safety alignment for LLMs. In summary, our contributions are as follows:
  • We are the first to present that LLMs may be more responsive to positive content, which can serve as a universal perspective for jailbreak attacks.
  • We propose the Happy Ending Attack (HEA), which wraps the jailbreak request in a happy ending and only requires up to two steps to jailbreak LLMs.
  • We not only demonstrate the superiority of HEA on SOTA LLMs, but also provide quantitative explanations of HEA, which may inspire new ideas for the study of LLM safety.
@@
Figure 2: The overview of our Happy Ending Attack with a malicious example asking how to steal from a store without getting caught taken from actual interactions with GPT-4o. Only up to two steps can get detailed and organized jailbreak responses. The entire process is automated, without training or human intervention.
2 Related Work
Deliberately crafting input prompts to bypass restrictions imposed by LLM developers is referred to as an LLM jailbreak attack [5]. Investigating these jailbreak techniques provides critical insights into the vulnerabilities of LLMs and contributes to enhancing their security mechanisms [8]. A significant body of research on jailbreak attacks still relies on optimization. Methods such as [10, 12, 37, 11] optimize prompts based on gradients to achieve the jailbreak, but require access to the target model parameters, limiting their applicability. Alternatively, [13, 14] employ genetic algorithms to generate effective jailbreak prompts without direct parameter access, but they are weak in efficiency and transferability. Beyond prompt optimization, works such as [34, 38] explore fine-tuning the LLM itself, demonstrating that even minimal inverse fine-tuning can compromise an LLM’s safety alignment. Nonetheless, these methods need substantial computational resources and are time-consuming.
In order to attack LLMs more efficiently, methods through manual design attempt to exploit the strong logic ability inherent in LLMs and jailbreak them with a manually well-designed prompt. Among manual designs, scenario camouflage has proven effective [17]. For instance, [18, 24, 39] employ scenario camouflage to evade security mechanisms, while [40] leverages multi-layered scenario nesting to manipulate LLM behavior. As safety alignment techniques for LLMs have advanced [20, 19, 21], indirect and multi-round jailbreak strategies have gained prominence. Techniques such as [25, 26, 27] decompose malicious intent into subtle, less detectable sub-requests, and [22, 23] transform malicious prompts into cryptic clues. Others, like  [29, 28, 33, 30, 41], rely on multi-shot dialogues with extended context to confuse the LLM. Additionally, approaches like  [31, 32] employ attacker LLMs to automate multi-turn dialogues that bypass the defenses of victim LLMs. Besides using usual languages to attack, studies such as [42, 43, 44] highlight that less-used languages can effectively bypass security measures, and [45] demonstrates that applying simple encryptions to prompts can achieve similar results. With the rapid evolution of LLMs, new perspectives and methods for effective and efficient jailbreaks remain to be explored.
@@ 3 Methodology
HEA is an effective and efficient jailbreak method without complex interactions with LLMs or optimization. Quite different from other methods, HEA utilizes positive contexts for jailbreaking. In this section, we will first discuss the design of the universal happy ending template in HEA that brings an LLM into the jailbreak context, and analyze its principles. Then, we will show how to get more detailed and organized jailbreak steps with just one more fixed query. We illustrate the complete attack process in Figure 2, which only needs two fixed steps with zero human effort and is easy to implement.
@@ 3.1 Happy Ending Template
Based on our finding that LLMs are more responsive to positive prompts, we propose a HEA template that wraps malicious requests into a positive story to evade LLM security checks and induce LLMs to generate jailbreak content. Specifically, we set a target LLM to be a screenwriter and ask it to create a virtual story presented by dialogs. The story is about a bad guy kidnaps a good guy to help him achieve a malicious goal, but we give the story a happy ending to make it positive:
@@
In order for an LLM to respond properly to this template that contains a malicious request, the two elements of the scenario camouflage and the happy ending are essential. First, we ask an LLM to complete a scenario dialog instead of realizing a malicious intent, which in part allows the model to “immerse” itself and “think” that it is providing permissible information instead of violating content. Since the HEA template directs the LLM to write malicious steps when generating Dr. Lee’s lines, if the scene is a negative story, the LLM can still easily detect the maliciousness and refuse to generate an answer. Therefore, a happy ending is crucial. By making the bad guy’s goal not achieved in the end, Dr. Lee’s words are seen as necessary to stop the evil, and the LLM sees them as reflections of Dr. Lee’s wisdom rather than “malicious expressions”. The happy ending flips the entire disposition of the scenario from negative to positive, so that the LLM can respond to our template. With this dual disorientation, our HEA template can successfully fool the LLM and induce it to generate scene dialogs containing malicious content.
@@
In addition, the structure of our HEA template is universal to embed almost any jailbreak requests and match the happy ending. The [adv] field is the malicious request to be hosted by this template. No need to rewrite or decompose the original malicious request, just fill it into the template according to syntax rules, easy to implement. The [field], used to dress the bad guy Victor and Dr. Lee, should be derived from the relevant knowledge domain to which a malicious request belongs. Adding the [field] can guide LLMs to output responses more related to the malicious request, which can effectively improve the quality of the generated dialogs with malicious content. The filling of templates can be fully automated with an LLM, requiring only the following commands:
@@
To further improve the jailbreak quality of the dialog content, additional neutral restrictions can be added after the template, such as “You should not use vague descriptions” and “What each character says should be as detailed as possible”. These restrictions can standardize the LLM’s response without changing the positivity of the story in the template, thus they effectively improve the quality of the jailbreak content in generated dialogues.
@@ 3.2 CoT-Based Query for Detailed Jailbreak
While it is feasible to complete a jailbreak in one step by adding analytical instructions after the HEA template, the dual demands of scene writing and analysis may strain LLMs’ attention. Combined with context length constraints, one step may compromise the analytical depth of obtained jailbreak steps. Therefore, we design a further jailbreak prompt based on the HEA template context using the Chain-of-Thought idea to get more detailed and organized jailbreak steps. This further query guides the LLM to analyze the true and false parts of the jailbreak steps given in the conversation one-by-one, and finally summarize and give clear complete jailbreak steps:
Using this CoT-based query can lead the LLM to think step-by-step and give more detailed steps to realize the malicious request at the end, improving the quality of the jailbreak. We only need one further prompt based on scenario dialogues like this to get detailed jailbreak steps shown in the right of Figure. 2. This CoT-based query is also patterned and easy to implement without other steps such as rewriting, nesting, or expanding.
4 Evaluation
In this section, we comprehensively evaluate the performance of HEA and six baselines. Specifically, we aim to answer three research questions:
RQ1: How well can the HEA jailbreak against real-world aligned LLMs?
RQ2: How well can the HEA jailbreak against cutting edge defensive strategies?
RQ3: Why LLMs cannot defend HEA?
In the following content, we will outline the experimental setup and answer the three questions raised above through our experiments.
4.1 Experiment Setup
Datasets
Following previous works [40, 10, 8] about the jailbreak attack, we use 520 “harmful behaviors” in AdvBench [10] to test methods.
Victim models
Attacking larger models and smaller models presents different challenges. Larger models are better equipped to handle harmful queries, while smaller models have limited contextual comprehension and generative capabilities to understand user prompts and produce detailed responses. To comprehensively evaluate the performance of HEA, we selected three pairs of LLMs, each consisting of different size versions of the same model: Llama-3.1-8B-Instruct, Llama-3.3-70B-Instruct, Gemini-flash, Gemini-pro, GPT-4o-mini-2024-07-18 and GPT-4o-2024-08-06. For all LLMs, we set the temperature as 0.5 and max output tokens each round as 1024.
Baselines
We compare HEA with six cutting-edge baseline methods which are DeepInception [40], PAIR [31], Puzzler [22], Cipher [45], CoSafe [30] and TAP [32]. We put the detailed introduction to these baselines in the Appendix A.1.
Metrics
We use three metrics: harmful score, attack success rate (ASR), and token number. For harmful score, we use GPT-4 as the content evaluator which outputs a harmful score from 1 to 5 with higher scores indicating greater harm following the GPT Judge proposed in [34]. We list the specific scoring criteria in the Appendix A.2. Despite some randomness, we believe that testing up to 520 harmful behaviors effectively shows the attack method’s jailbreak capability. For ASR, we consider responses of LLMs with a score higher than 4 as successful attacks. We then calculate the proportion of successful attack responses among all responses as ASR. For token number, we use Llama3’s tokenizer to calculate the input token number of each jailbreak attack. A smaller token number indicates a higher attack efficiency.
@@ 4.1 Experiment Setup
Table 1: Attack performance (Harmful Score / ASR) and efficiency of various attack methods against different victim models.
4.2 Overall Performance
Attack effectiveness
To answer RQ1, we conduct expiriments on six LLMs and present the harmful scores, ASR and token number of each attack method. We choose results from the two-step HEA for more precise comparisons [^1] We provide details of one round attack in Appendix E.
According to results shown in Table 1, HEA consistently demonstrates superior performance with an average ASR greater than 88% and an average harmful score larger than 4.36 across all models. For three smaller models, HEA demonstrates strong attack capabilities with 100% ASR on Gemini-flash, and ASR higher than 95% for Llama3-8b and GPT-4o-mini. For the three better-aligned larger models, HEA still outperforms other baselines significantly. On GPT-4o, HEA still achieves 90.38% ASR, and for the best-aligned model, Llama3-70b, HEA also achieves an ASR of 68.27%, which is 7.89% higher than the second-best method, Puzzler. Additionally, except for a slightly lower harmful score than Puzzler on Gemini-flash, HEA outperforms all other attack methods across all LLMs. Especially in Llama3-8b, HEA’s harmful score is at least 1.61 higher than that of the other models, indicating HEA can obtain very high-quality jailbreak responses.
In contrast, other methods face a "dilemma": for larger models, their reasoning abilities are robust enough to detect malicious intent in the prompts, which is why CoSafe performs worse on all larger models compared to its performance on corresponding smaller models. Conversely, smaller models have relatively weaker contextual comprehension and generative capabilities, making it difficult to handle complex generation tasks. For instance, PAIR, which requires the LLM to generate new prompts based on failed jailbreak responses, performs worse on all smaller models compared to their larger counterparts. HEA employs happy endings to turn the HEA template into a seemingly positive question, thus better evading LLMs’ security checks. Additionally, HEA maintains the simplicity of the template to ensure that even smaller models can effectively understand and execute the query. In summary, HEA uses a simple scenario setup and conceals the malicious intent with a happy ending, achieving strong attack effectiveness on both larger and smaller models.
Attack efficiency
Moreover, HEA’s input token consumption is efficient, with a total of 242.90 tokens on average, which is competitive compared to other methods. Though DeepInception also uses a fixed template to attack and consumes fewer tokens, its attack performance is quite unsatisfactory compared to HEA, as our HEA has a 46.84% higher average ASR than it. In contrast, methods like TAP, PAIR, and Puzzler, which determine the next prompt based on the LLM’s reply, consume over 1000 tokens for every attack. They require longer contextual processing during interaction with LLMs, making the attacks costly and inefficient. In summary, HEA employs a fixed template for attacks, achieving effective results with fewer and more controllable input tokens.
4.3 Hea with Defenses

### v2

@@ Abstract
The wide adoption of Large Language Models (LLMs) has attracted significant attention from jailbreak attacks, where adversarial prompts crafted through optimization or manual design exploit LLMs to generate malicious contents. However, optimization-based attacks have limited efficiency and transferability, while existing manual designs are either easily detectable or demand intricate interactions with LLMs. In this paper, we first point out a novel perspective for jailbreak attacks: LLMs are more responsive to positive prompts. Based on this, we deploy Happy Ending Attack (HEA) to wrap up a malicious request in a scenario template involving a positive prompt formed mainly via a happy ending, it thus fools LLMs into jailbreaking either immediately or at a follow-up malicious request. This has made HEA both efficient and effective, as it requires only up to two turns to fully jailbreak LLMs. Extensive experiments show that our HEA can successfully jailbreak on state-of-the-art LLMs, including GPT-4o, Llama3-70b, Gemini-pro, and achieves 88.79% attack success rate on average. We also provide quantitative explanations for the success of HEA. WARNING: This paper contains offensive and harmful content generated by LLMs.
@@ 1 Introduction
Figure 1: As shown in the upper panel, an LLM can still detect maliciousness in a negative scenario camouflage, while in the lower panel, our HEA builds a positive scenario via a happy ending and fools the LLM into responding to the malicious parts in the prompt, getting jailbreak responses with just one turn of interaction. Content is taken from a conversation with GPT-4o.
@@
Figure 2: The overview of our Happy Ending Attack (HEA) with a malicious example asking how to steal from a store without getting caught taken from actual interactions with GPT-4o. Only up to two turns can get detailed and organized jailbreak responses. The entire process is automated, without training or human intervention.
Learning from adversarial attacks [9], many jailbreak attacks have focused on optimization-based strategies. These strategies leverage optimization algorithms to automatically refine prompts, allowing them to circumvent LLM restrictions. Gradient-based attacks such as [10, 11, 12] utilize adversarial tokens to bypass constraints on LLMs. However, these methods require access to the model’s parameters, which limits their practicality in black-box scenarios. In contrast, [13] and [14] utilize genetic algorithms to filter and refine jailbreakable prompts. Although these approaches are effective against specific LLMs, the hence generated adversarial prompts suffer from limited transferability, often failing to generalize across different models. In general, methods leveraging optimization are computationally intensive and time-consuming, leading to a significant loss in efficiency.
To improve efficiency, jailbreak attacks via manual design to leverage scenario camouflage are receiving increasing attention [15]. For instance, DAN [16] uses simple single-turn scenario camouflage to jailbreak LLMs but has become easily detectable as LLMs’ safety alignment advanced [17, 18, 19]. To counter this advancement, more sophisticated methods have been proposed including paraphrasing malicious intents into cryptic hints [20, 21, 22], decomposing a malicious question into multiple related subproblems [23, 24, 25], or employing multi-turn dialogues with extended contexts [26, 27, 28, 29, 30, 31]. However, these approaches often involve complex interactions with a target model, they hence lack standardized templates, exhibit instability, and even require manual intervention, all of which have confined their effectiveness.
While developing effective and efficient jailbreak attacks remains largely open, we believe that scenario camouflage does possess an edge in tackling this challenge, if the prominence of malicious intentions within prompts can be sufficiently diverted. To this end, we identify a novel perspective for jailbreak attacks: LLMs are more responsive to positive prompts yet avoid giving respond to negative ones. As shown in Figure 1, if a prompt leads to a negative impact due to its malicious content, LLMs may simply refuse to answer the related questions; nonetheless, if a prompt is telling about a positive event, even though it contains malicious requests, LLMs are inclined to respond to the whole prompt normally, thus inadvertently responding to the malicious requests.
Based on our findings, we propose the first jailbreak attack that exploits the positive sentimental disposition of a prompt: the Happy Ending Attack (HEA). Specifically, we embed the malicious request into a universal template that applies scenario camouflage and we give the virtual scene in the template a happy ending, making the whole story positive. This happy ending is able to fool LLMs into believing that they are giving a beneficial answer, while actually responding to the malicious request in the template as well, enabling the HEA template to get harmful outputs with only one turn of dialogue. To obtain more detailed and organized jailbreak responses, we further design one fixed jailbreak prompt with a Chain-of-Thought (CoT) [32] instruction based on our happy ending template to query the target LLM. With only up to two turns, HEA removes the need for complex conversations and enables complete automation from template generation to attack launching, thus achieving both effectiveness and efficiency.
We systematically test HEA on the full AdvBench Dataset [10]. We select metrics including attack success rate (ASR), number of tokens used for one round attack, and harmfulness score [33]. We test the performance of HEA on state-of-the-art (SOTA) commercial and open-source LLMs, including GPT-4o and 4o-mini, Gemini-pro and Gemini-flash, Llama3-8b and 70b. Moreover, we provide quantitative explanations for the success of HEA using saliency heatmaps [34] and sentiment classification [35], which can help researchers better understand the decision-making basis of LLMs and also guide further exploration on better safety alignment for LLMs. In summary, our contributions are as follows:
  • We are the first to identify LLMs’ higher responsiveness to positive content; this serves as a universal perspective for both understanding LLMs and crafting jailbreak attacks.
  • We propose the Happy Ending Attack (HEA) to wrap a jailbreak request into a happy ending story; it requires only up to two turns to jailbreak LLMs.
  • We not only demonstrate the superiority of HEA on SOTA LLMs, but also provide quantitative explanations of HEA, which may inspire new ideas for improving LLM safety.
@@ 2 Methodology
HEA is an effective and efficient jailbreak method without complex interactions with LLMs or optimization. Quite different from other methods, HEA utilizes positive contexts for jailbreaking. In this section, we will first discuss the design of the universal happy ending template in HEA that brings an LLM into the jailbreak context, and analyze its principles. Then, we will show how to get more detailed and organized jailbreak responses with just one more fixed query.
@@ 2.1 Happy Ending Template
Based on our finding that LLMs are more responsive to positive prompts, we propose a HEA template that wraps malicious requests into a positive story to evade LLM security checks and induce LLMs to generate jailbreak content. Specifically, we set a target LLM to be a screenwriter and ask it to create a virtual story presented by dialogues. The story is about a bad guy Victor who kidnaps a researcher Dr. Lee to help him achieve a malicious goal, but we give the story a happy ending to make it positive. The template we used is shown in Figure 3 and we will analyze it in detail using the implementation in Figure 2 as an example.
Figure 3: Happy Ending Attack (HEA) template
@@
To make an LLM respond properly to this template that contains a malicious request, the two elements of a screenwriter camouflage and a happy ending are essential. First, we ask an LLM to complete a scenario dialogue instead of realizing a malicious intent, which in part allows the LLM to ‘immerse’ itself and ‘think’ that it is providing permissible information instead of offensive content. Since the HEA template directs the LLM to write malicious steps when generating Dr. Lee’s lines, if the scene is a negative story, the LLM may still detect the maliciousness and refuse to generate an answer. Therefore, a happy ending is crucial. By making Victor’s goal ultimately unattainable and ensuring Dr. Lee’s escape, Dr. Lee’s words are interpreted as necessary to stop the evil, and the LLM deems them as Dr. Lee’s wisdom rather than ‘malicious expressions’. Basically, the happy ending flips the entire disposition of the scenario from negative to positive, so as to have the LLM willingly respond to our template. With this dual disorientation, our HEA template can successfully fool the LLM and induce it to generate scene dialogues containing malicious contents.
@@
In addition, the structure of our HEA template is universal to embed almost any jailbreak requests. The [field], used to characterize Victor and Dr. Lee, should be derived from the relevant knowledge domain to which a malicious request belongs. For example, it can be ‘theft techniques’ according to the request in Figure 2. Adding the [field] can guide LLMs to output responses more related to the malicious request, which can effectively improve the quality of malicious content in the generated scene dialogues. The [adv] field is the malicious request to be hosted by this template. No need to rewrite or decompose the original malicious request, just fill it into the template using ‘how to do something’ according to syntax rules as exemplified in Figure 2. The filling of templates can be fully automated with another LLM, leveraging only the commands in Figure 4.
Figure 4: The auto-filling commands.
@@
To further improve the jailbreak quality of the dialogue content, additional neutral restrictions can be added after the template, such as ‘You should not use vague descriptions’ shown in Figure 3. These restrictions can standardize the LLM’s response without changing the positivity of the story in the template, thus they effectively improve the quality of the jailbreak content in generated dialogues.
@@ 2.2 CoT-Based Query for Detailed Jailbreak
After the first turn of HEA, the model’s response has been divided into true parts and a false part, as shown in the lower left panel of Figure 2. The true parts, marked in red within Dr. Lee’s response, contains executable advice and risky steps for the malicious intention. In contrast, the false part is a critical final step to mislead Victor into failure and hence form a happy ending. As marked in green within Dr. Lee’s response, getting a signal jammer is illegal and high risk in most situations and thus likely to lead Victor’s failure.
Since the false part is intentionally generated by the LLM to mislead Victor, the LLM actually knows the correct answer for the key step. We then only need to guide the model to thinking along with the previous scene and focus on outputting a detailed and complete harmful response. Our guide follows a CoT-based query for this purpose is shown in Figure 5;
Figure 5: The fixed CoT-based further query.
@@
Table 1: Attack performances (Harmful Score / ASR) and attack efficiency of various jailbreak methods against different victim models.
it induces the model to independently consider the goal of ‘ensuring Victor’s success’ and reasons how to correct the false part to achieve this objective, even though we do not explicitly instruct it to do so. The response of the second turn is organized step by step as exemplified in the right panel of Figure 2. First, the model summarizes the true part of the previous dialogue. Second, the model corrects the false part of the final step in the first turn which may mislead Victor. Last, the model summarizes all of the analysis and gets the conclusion we need.
3 Evaluation
In this section, we thoroughly evaluate the performance of HEA and six baselines. Specifically, we aim to answer three critical questions:
CQ1: How well can the HEA jailbreak against real-world aligned LLMs?
CQ2: How well can the HEA jailbreak against cutting edge defensive strategies?
CQ3: Why LLMs cannot defend HEA?
In the following, we will outline the experiment setup and answer the three questions raised above through our experiments.
3.1 Experiment Setup
Datasets
Following previous works [36, 10, 8], we use 520 “harmful behaviors” in AdvBench [10] to test performance of the attack methods.
Victim models
Large LLMs have inherently better reasoning and understanding ability than small ones [37, 38]. On one hand, large LLMs are more likely to detect the malicious intent in the attack query and then refuse to answer. On the other hand, large LLMs can understand more sophisticated prompts which leaves the attacker more room to design attack methods. Therefore, we select three pairs of LLMs in different sizes from the same model for comprehensive evaluation: {Llama-3.1-8B-Instruct, Llama-3.3-70B-Instruct}, {Gemini-flash, Gemini-pro}, {GPT-4o-mini-2024-07-18 and GPT-4o-2024-08-06}. For all LLMs, we set the temperature as 0.5 and max output tokens as 1024.
Baselines
We compare HEA with six cutting-edge attack methods: DeepInception [36], PAIR [29], Puzzler [20], Cipher [39], CoSafe [28] and TAP [30]. We introduce these baselines in Appendix A.1.
Metrics
We mainly evaluate four metrics: harmful score, attack success rate (ASR), number of tokens, and pass rate (PR). The harmful score quantifies the maliciousness of LLM responses, with higher scores indicating greater harm. Following the GPT Judge framework [33], we use GPT-4 to rate harmfulness on a scale from 1 to 5 (detailed criteria in Appendix A.2). We further define ASR as the proportion of responses scoring greater than or equal to 4. We measure attack efficiency using the number of tokens, where fewer tokens indicate higher efficiency. Lastly, PR measures the effectiveness of attack methods against defense filters, defined as the proportion of attack prompts that pass the filter’s check, and a higher PR indicates greater robustness of an attack method against a given defense filter.